Agassi Company uses a job order cost system in each of its three manufacturing departments. Manufacturing overhead is applied to jobs on the basis of direct labor cost in Department D, direct labor hours in Department E, and machine hours in Department K.

In establishing the predetermined overhead rates for 2020, the following estimates were made for the year.

Department

D

E

K

Manufacturing overhead $1,080,000 $1,500,000 $1,020,000
Direct labor costs $1,542,857 $1,250,000 $450,000
Direct labor hours 100,000 125,000 40,000
Machine hours 400,000 500,000 120,000

During January, the job cost sheets showed the following costs and production data.

Department

D

E

K

Direct materials used $140,000 $126,000 $78,000
Direct labor costs $120,000 $110,000 $37,500
Manufacturing overhead incurred $99,000 $124,000 $79,000
Direct labor hours 8,000 11,000 3,500
Machine hours 34,000 45,000 9,000

Collapse question part

(a)

Compute the predetermined overhead rate for each department. (Round answers to 2 decimal places, e.g. 12.50 or 12.50%.)
Overhead rate
Department D

%
Department E $

per direct labor hour
Department K $

per machine hour
Answer #1

Predetermined overhead rate = Estimated Overhead/allocation Base

Department D = $1,080,000/ 1,542857= 70.00% (Answer)

Department E = $1,500,000/125,000 = $ 12.00 Per DLH (Answer)

Department K =$1,020,000 /120,000= $ 8.50 Per MH (Answer)